OpenStudy (anonymous):

Multiply (9x+1)(2x^2+9x+2) i got 18x^3+83x^2+27x+2

OpenStudy (lgbasallote):

that's correct

OpenStudy (anonymous):

1) (9x + 1)(x + 2) = 9x(x) + 9x(2) + 1(x) + 1(2) = 9x^2 + 18x + x + 2 = 9x^2 + 19x + 2

OpenStudy (anonymous):

u got it ;)